# Stochastic block models with many communities and the Kesten–Stigum bound

### 04 Mar 2025 v1

@@ Abstract
We study the inference of communities in stochastic block models with a growing number of communities. For block models with $n$ vertices and a fixed number of communities $q$, it was predicted in Decelle et al. (2011) that there are computationally efficient algorithms for recovering the communities above the Kesten–Stigum (KS) bound and that efficient recovery is impossible below the KS bound. This conjecture has since stimulated a lot of interest, with the achievability side proven in a line of research that culminated in the work of Abbe and Sandon (2018). Conversely, recent work by Sohn and Wein (2025) provides evidence for the hardness part using the low-degree paradigm.
@@
Our low-degree hardness results also naturally have consequences for graphon estimation, improving results of Luo and Gao (2024).
@@ 1.1 Main results
Here, $d$ is the average degree of the block model and $\lambda$ is a measure of signal strength. With this parametrization, the relevant regime for the study of recovery is when $d,\lambda=\Theta(1)$. In our original parameters, this translates to $a=\Theta(q)$ and $b=\Theta(1)$. Recall that the Kesten–Stigum (KS) threshold is $d\lambda^{2}=1$.
@@ 1.1.2 Low-degree hardness of recovery
Here, all expectations are with respect to the joint distribution of $(x,Y)$ drawn from the block model $\mathsf{SBM}(n,q,d,\lambda)$. The normalization guarantees that $\mathsf{Corr}_{\leq D}\in[0,1]$, where $\mathsf{Corr}_{\leq D}=1-o(1)$ indicates near-perfect recovery of $x$, while $\mathsf{Corr}_{\leq D}=\Omega(1)$ indicates non-trivial correlation with $x$, and $\mathsf{Corr}_{\leq D}=o(1)$ indicates failure to correlate with $x$. It is a standard fact (see e.g. [52, Fact 1.1]) that $\mathsf{Corr}_{\leq D}$ is directly related to the minimum mean squared error in estimating $\mathds{1}\{\sigma^{\star}_{1}=\sigma^{\star}_{2}\}$ by a degree-$D$ polynomial.
Bounds on $\mathsf{Corr}_{\leq D}$ have a conjectural interpretation for time complexity: if $\mathsf{Corr}_{\leq D}=o(1)$ for some super-logarithmic degree, i.e. $D\geq(\log n)^{1+\Omega(1)}$, this suggests that any algorithm needs super-polynomial time to obtain non-trivial correlation with $x$ (see [27, Conjecture 2.2.4]). Furthermore, if $\mathsf{Corr}_{\leq n^{c}}=o(1)$ for a constant $c>0$, this suggests that runtime $\exp(n^{c-o(1)})$ is required (which is roughly the runtime needed to evaluate a degree-$n^{c}$ polynomial term-by-term). These rules are justified by their success in tracking the best known time complexity for a variety of different statistical models, along with the fact that degree-$O(\log n)$ polynomials tend to capture some of our best algorithmic tools like spectral methods. We note that there are limits to the style of models where the above conjectures hold up, but the SBM is considered a “core” problem in this area, where low-degree polynomials have already been “stress-tested.” In particular, for fixed $q$, the low-degree threshold coincides with the KS threshold for both hypothesis testing [29, 4] and weak recovery [55, 17]. The techniques of [55] extend to $q\ll n^{1/8}$. A different approach relies on a low-degree hardness conjecture to establish average-case hardness of weak recovery for $q=n^{o(1)}$ [17].
Our hardness results show that if $q\ll n^{1/2}$ then polynomials of degree $n^{c}$ fail to achieve non-trivial correlation with $\mathds{1}\{\sigma^{\star}_{1}=\sigma^{\star}_{2}\}$ below the KS bound.
@@
We emphasize that the conditions $q=n^{1/2-\Omega(1)}$, and $d\lambda^{2}\leq 1$ in Theorem 1.4 are necessary in the sense that if either one of them fails, then $\mathsf{Corr}_{\leq O(\log n)}\geq\delta$. We refer to Corollary 1.5 and the following remark.
@@ 1.1.4 Computational barriers to graphon estimation
Graphon estimation has gained great interest in the last decade [57, 3, 9, 22, 32]. The goal is to estimate the graphon $f$ given the adjacency matrix $Y$. In particular, it was shown in [33, Section 3] that estimating $f$ in squared error can be achieved by estimating $M$ in Frobenius norm:
@@ 2.1 Algorithm preliminaries
With high probability, the set $U_{*}$ consists of vertices of degree at least $\frac{1-\chi}{2}\frac{\log n}{\log\log n}$, and moreover there are at least $\frac{q}{4}$ communities with a unique representative in $U_{*}$. Moreover, among all neighbors of $U_{*}$, no uniquely represented community appears more than $\log n$ times.
@@
Uniformly choose a subset $U\subseteq V$ of size $n^{\frac{\chi+1}{2}}$. The degree distribution of each vertex to $V\setminus U$ is an independent $\mathrm{Bin}((1-o(1))n,\frac{d}{n})$ random variable. By a Poisson approximation, the probability that a vertex has degree at least $t$ is
$\mathbb{P}\left(\mathrm{Bin}\left((1-o(1))n,\frac{d}{n}\right)>t\right)\geq\frac{d^{t}}{t!}e^{-d}\approx\exp(t\log d-d-t\log t-t).$
When $t=\frac{1-\chi}{2}\cdot\frac{\log n}{\log\log n}$, this probability is roughly $n^{\frac{\chi-1}{2}}$. Thus, for large enough $n$, with high probability we have at least $n^{\frac{\chi-1}{2}}n^{\frac{\chi+1}{2}}=q$ vertices in $U$ of degree at least $\frac{1-\chi}{2}\cdot\frac{\log n}{\log\log n}$. This confirms the first claim about $U_{*}$.
For the second claim, let $X$ be the number of communities with a unique representative in $U_{*}$. Then $\mathbb{E}\left[X\right]=q\left(\frac{q-1}{q}\right)^{q-1}=(1+o(1))\frac{q}{e}$ and
$\text{Var}\left(X\right)$ $=\mathbb{E}\left[X^{2}\right]-\mathbb{E}\left[X\right]^{2}=(1+o(1))\frac{q}{e}+(q-1)^{2}\left(\frac{q-2}{q}\right)^{q-2}-(1+o(1))\frac{q^{2}}{e^{2}}=o(q^{2}).$
Thus, by Chebyshev’s inequality we deduce that $X\geq\frac{q}{4}$ with high probability.
For the final claim, observe that the number of occurrences of each uniquely represented community is dominated by $\mathrm{Bin}(\frac{1-\chi}{2}\frac{\log n}{\log\log n},\frac{1+(q-1)\lambda}{q})+\mathrm{Bin}(q\frac{1-\chi}{2}\frac{\log n}{\log\log n},\frac{1-\lambda}{q})$. In particular, it has expectation on the order of $\frac{\log n}{\log\log n}$. Standard concentration of binomial random variables implies the desired result. ∎
@@ 2.2 Path counting
In this subsection, we recall the combinatorial path bounds from [45, Section 4] that control the number of paths of fixed length in the complete graph with conditions on the number of self-intersections. These will be a crucial ingredient in controlling the moments of the weighted path statistics in the next section. We define a path to be a sequence of vertices in the complete graph where consecutive vertices are distinct. A path is called self-avoiding if all of its vertices are distinct.
@@ 2.3 Weighted path statistics
2 Let the $q$ vertices of highest degree be $U_{*}=\{u_{1},\ldots,u_{q}\}$.
@@
2 Let the $q$ vertices of highest degree be $U_{*}=\{u_{1},\ldots,u_{q}\}$.
@@ 3.3 Construction of $u$
We construct $u$ in the same manner as in [55, Section 6.3]. Namely, it is constructed implicitly by the following recursion:
@@ 4 Information-theoretic threshold
Part 2 of the theorem — the impossibility result — follows from a percolation style argument. We first show that we may reduce to the case where only the within-community edges are revealed to us. Consider generating a sample from $G\sim\mathsf{SBM}(n,q,d,\lambda)$ as follows. Assign each vertex $v$ an i.i.d. uniform $\sigma^{\star}_{v}\in[q]$ community label as before. For each pair of vertices $u,v$ such that $\sigma^{\star}_{u}=\sigma^{\star}_{v}$ connect them with an edge independently with probability $\frac{qd\lambda}{n}\cdot\frac{1}{1-\frac{d(1-\lambda)}{n}}=(1+o(1))\frac{qd\lambda}{n}$ to obtain a graph $\widetilde{G}$. Finally, let $G=\widetilde{G}\cup\widehat{G}$ where $\widehat{G}$ is the Erdős–Rényi graph $G\left(n,\frac{d(1-\lambda)}{n}\right)$. We claim that $\delta$-recovery is possible in $\mathsf{SBM}(n,q,d,\lambda)$ only if recovery is possible using $\widetilde{G}$. Indeed, given a sample of $\widetilde{G}$ we can simply sample $\widehat{G}$ and perform recovery from the block model. Thus, it suffices to show that $\delta$-recovery is impossible from $\widetilde{G}$.

### 11 Jun 2025 v2

@@ Abstract
We study the inference of communities in stochastic block models with a growing number of communities. For block models with $n$ vertices and a fixed number of communities $q$, it was predicted in Decelle et al. (2011) that there are computationally efficient algorithms for recovering the communities above the Kesten–Stigum (KS) bound and that efficient recovery is impossible below the KS bound. This conjecture has since stimulated a lot of interest, with the achievability side proven in a line of research that culminated in the work of Abbe and Sandon (2018). Conversely, recent work provides evidence for the hardness part using the low-degree paradigm.
@@
Our low-degree hardness results also naturally have consequences for graphon estimation, improving results of Luo and Gao (2024). [^1] Accepted for presentation at the Conference on Learning Theory (COLT) 2025
@@ 1.1 Main results
Here, $d$ is the average degree of the block model and $\lambda$ is a measure of signal strength. With this parametrization, the relevant regime for the study of recovery is when $d,\lambda=\Theta(1)$. Note that when $q=q_{n}\to\infty$, $\lambda=\Theta(1)$ implies that $\lambda\geq 0$ (assortative) since $\lambda\in[-(q-1)^{-1},1]$ by definition. In our original parameters, this translates to $a=\Theta(q)$ and $b=\Theta(1)$. This will be the scaling assumed for the rest of the paper, unless otherwise noted. Recall that the Kesten–Stigum (KS) threshold is $d\lambda^{2}=1$.
@@ 1.1.2 Low-degree hardness of recovery
Here, all expectations are with respect to the joint distribution of $(x,Y)$ drawn from the block model $\mathsf{SBM}(n,q,d,\lambda)$. The normalization guarantees that $\mathsf{Corr}_{\leq D}\in[0,1]$, where $\mathsf{Corr}_{\leq D}=1-o(1)$ indicates near-perfect recovery of $x$, while $\mathsf{Corr}_{\leq D}=\Omega(1)$ indicates non-trivial correlation with $x$, and $\mathsf{Corr}_{\leq D}=o(1)$ indicates failure to correlate with $x$. It is a standard fact (see e.g. [54, Fact 1.1]) that $\mathsf{Corr}_{\leq D}$ is directly related to the minimum mean squared error in estimating $\mathds{1}\{\sigma^{\star}_{1}=\sigma^{\star}_{2}\}$ by a degree-$D$ polynomial.
Bounds on $\mathsf{Corr}_{\leq D}$ have a conjectural interpretation for time complexity: if $\mathsf{Corr}_{\leq D}=o(1)$ for some super-logarithmic degree, i.e. $D\geq(\log n)^{1+\Omega(1)}$, this suggests that any algorithm needs super-polynomial time to obtain non-trivial correlation with $x$ (see [29, Conjecture 2.2.4]). Furthermore, if $\mathsf{Corr}_{\leq n^{c}}=o(1)$ for a constant $c>0$, this suggests that runtime $\exp(n^{c-o(1)})$ is required (which is roughly the runtime needed to evaluate a degree-$n^{c}$ polynomial term-by-term). These rules are justified by their success in tracking the best known time complexity for a variety of different statistical models, along with the fact that degree-$O(\log n)$ polynomials tend to capture some of our best algorithmic tools like spectral methods. We note that there are limits to the style of models where the above conjectures hold up, but the SBM is considered a “core” problem in this area, where low-degree polynomials have already been “stress-tested.” In particular, for fixed $q$, the low-degree threshold coincides with the KS threshold for both hypothesis testing [31, 4] and weak recovery [57, 17]. The techniques of [57] extend to $q\ll n^{1/8}$. A different approach relies on a low-degree hardness conjecture to establish average-case hardness of weak recovery for $q=n^{o(1)}$ [17]. Coarser low-degree lower bounds for recovery are given by [40], which apply for all $q\ll n$ but are not sharp enough to capture the precise KS transition.
Our hardness results show that if $q\ll n^{1/2}$ then polynomials of degree $n^{c}$ fail to achieve non-trivial correlation with $\mathds{1}\{\sigma^{\star}_{1}=\sigma^{\star}_{2}\}$ below the KS bound. Notice that this theorem holds beyond the sparse regime, and even with constant edge probabilities with a modified SNR.
@@
Neither of the conditions $q=n^{1/2-\Omega(1)}$ and $d\lambda^{2}\leq 1$ are removable, as explained in the remark following Corollary 1.5.
@@ 1.1.3 Phase transition at the KS bound
The low-degree threshold for $q\gg\sqrt{n}$ remains an open question, and we do not give a low-degree lower bound in this regime. The existing lower bound of [40] covers the full range of $q\ll n$, and when specialized to our setting where $b,d=\Theta(1)$ and $a,q\to\infty$, this shows low-degree hardness of recovery when $\lambda^{2}\ll\min\{1,n/q^{2}\}$ where $\ll$ hides $\log(n)$ factors. For $q\ll\sqrt{n}$ this matches the KS threshold up to log factors, and for $q\gg\sqrt{n}$ this gives a lower bound that is non-trivial but far from our positive result $d\lambda^{1/\chi}>C(\log d)^{2}$ since $\lambda$ needs to shrink with $n$.
@@ 1.1.4 Computational barriers to graphon estimation
Graphon estimation has gained great interest in the last decade [59, 3, 9, 23, 34]. The goal is to estimate the graphon $f$ given the adjacency matrix $Y$. In particular, it was shown in [35, Section 3] that estimating $f$ in squared error can be achieved by estimating $M$ in Frobenius norm:
@@ 1.3 Discussion and future directions
We note that some recent work [21] on clustering high-dimensional Gaussian mixture models has some parallels to ours. This problem shares some similarities to the SBM, where the number of communities $q$ corresponds to the number of clusters, and the number of vertices $n$ corresponds to the number of samples. The work [21] identifies the computational threshold at an analogous KS bound up to log factors, giving a low-degree lower bound that matches the known poly-time algorithms. Curiously, the behavior changes once $q$ exceeds $\sqrt{n}$, just like in the SBM. When $q\gg\sqrt{n}$, the optimal poly-time algorithm for clustering is a simple one based on grouping nearby points [21, Appendix F]. A non-rigorous transposition (supported by e.g., [26]) suggests a similar change in computational threshold for rather dense block models. However, this transposition breaks down for constant $d$ and it is not clear that a similarly simple algorithm works. Since we focus on the sparse setting of the SBM we leave this potential connection to future research.
@@ 2.1 Algorithm preliminaries
Suppose $q=o(\frac{n}{(\log\log n)^{\log\log n}})$. Let $U$ be a set of size $\sqrt{nq}$ and let $U_{*}$ be a uniformly random subset of size $q$ drawn from all vertices in $U$ which have degree at least $\log\log n$ to $V\setminus U$. With high probability, there are at least $\frac{q}{4}$ communities with a unique representative in $U_{*}$.
@@
We reveal $\sigma^{\star}$ so that all edges in the graph are drawn independently from each other. Let $U_{i}$ be the vertices in $U$ with label $i$, and $V_{i}$ be the vertices in $V\setminus U$ with label $i$. With high probability,
$|U_{i}|=\sqrt{\frac{n}{q}}+O\left(\left(\frac{n}{q}\right)^{1/4}\log(\frac{n}{q})\right)\text{ and }|V_{i}|=\frac{n}{q}+O\left(\sqrt{\frac{n}{q}}\log(\frac{n}{q})\right)\quad\forall i.$
The degrees of vertices in $U$ to $V\setminus U$ are independent, as we have revealed all of the community labels. Moreover, each vertex $u_{i}\in U_{i}$ has degree distribution
$D_{i}=\mathrm{Bin}\left(|V_{i}|,\frac{a}{n}\right)+\mathrm{Bin}\left(n-\sqrt{nq}-|V_{i}|,\frac{b}{n}\right)$
where $+$ denotes convolution in this context. This differs from
$\widetilde{D_{i}}=\mathrm{Pois}\left(|V_{i}|\cdot\frac{a}{n}+\left(n-\sqrt{nq}-|V_{i}|\right)\cdot\frac{b}{n}\right)=\mathrm{Pois}\left(d+O\left(\sqrt{\frac{q}{n}}\log(\frac{n}{q})\right)\right)$
by $O\left(\frac{q}{n}\right)$ in total variation distance (see e.g. Lemma 5 in [46]). We compare $\mathbb{P}\left(\widetilde{D_{i}}>\log\log n\right)$ with $\mathbb{P}\left(D>\log\log n\right)$ where $D\sim\mathrm{Pois}(d)$. First notice that
$\mathbb{P}\left(\widetilde{D}_{i}>\log\log n\right)\wedge\mathbb{P}\left(D>\log\log n\right)>\frac{1}{(\log\log n)^{\log\log n}}$
A standard Chernoff bound yields
$\mathbb{P}\left(\widetilde{D_{i}}>\log n\right)\vee\mathbb{P}\left(D>\log n\right)<\frac{(ed)^{\log n}e^{-d}}{(\log n)^{\log n}}=o\left(\frac{1}{(\log\log n)^{\log\log n}}\right).$
Thus we have
$\mathbb{P}\left(D>\log\log n\right)=(1+o(1))\sum_{k=\log\log n}^{\log n}e^{-d}\frac{d^{k}}{k!}$
and
$\mathbb{P}\left(\widetilde{D}_{i}>\log\log n\right)=(1+o(1))\sum_{k=\log\log n}^{\log n}e^{-d-O(\sqrt{q/n}\log(n/q))}\frac{(d+O(\sqrt{q/n}\log(n/q)))^{k}}{k!}.$
For each $\log\log n<k<\log n$, we have
$\frac{e^{-d}\frac{d^{k}}{k!}}{e^{-d-O(\sqrt{q/n}\log(n/q))}\frac{(d+O(\sqrt{q/n}\log(n/q)))^{k}}{k!}}$ $=e^{O(\sqrt{q/n}\log(n/q))}\left(1+O\left(\sqrt{\frac{q}{n}}\log(n/q)\right)\right)^{k}$
$\leq\exp(O\left(k\sqrt{\frac{q}{n}}\log(\frac{n}{q})\right))=1+o(1).$
Thus, all together we have shown that $\mathbb{P}\left(\widetilde{D}_{i}>\log\log n\right)=(1+o(1))\mathbb{P}\left(D>\log\log n\right)$ for every $i$. Now, the number of vertices in $U$ with label $i$ and degree larger than $\log\log n$ is distributed as
$\mathrm{Bin}\left(|U_{i}|,\mathbb{P}\left(D_{i}>\log\log n\right)\right).$
Thus, with high probability, there are
$|U_{i}|\cdot\mathbb{P}\left(D_{i}>\log\log n\right)=(1+o(1))\sqrt{\frac{n}{q}}\cdot\mathbb{P}\left(D>\log\log n\right)$
such vertices. Here we used that the expectation tends to infinity. Notice that the right hand side is independent from $i$. Thus, a randomly selected vertex of $U$ with degree at least $\log\log n$ has label $i$ with probability $\frac{1+o(1)}{q}$ for every $i$. If we let $X$ denote the number of communities with a unique representative selected in $U_{*}$, then
$\mathbb{E}\left[X\right]=q^{2}\cdot\frac{1+o(1)}{q}\cdot\left(1-\frac{1+o(1)}{q}\right)^{q-1}=(1+o(1))\frac{q}{e}$
and
$\text{Var}\left(X\right)=(1+o(1))\frac{q}{e}+q^{2}(q-1)^{2}\left(\frac{1+o(1)}{q}\right)^{2}\left(1-\frac{1+o(1)}{q}\right)^{q-2}-(1+o(1))\frac{q^{2}}{e^{2}}=o(q^{2}).$
By Chebyshev’s inequality we deduce that $X\geq\frac{q}{4}$ with high probability. ∎
@@ 2.2 Path counting
In this subsection, we recall the combinatorial path bounds from [47, Section 4] that control the number of paths of fixed length in the complete graph with conditions on the number of self-intersections. These will be a crucial ingredient in controlling the moments of the weighted path statistics in the next section. We define a path to be a sequence of vertices in the complete graph where consecutive vertices are distinct. A path is called self-avoiding if all of its vertices are distinct.
@@ 2.3 Weighted path statistics
2 Draw $q$ vertices $U_{*}=\{u_{1},\ldots,u_{q}\}$ uniformly at random from all vertices in $U$ having degree at least $\log\log n$ to $V\setminus U$.
@@
2 Draw $q$ vertices $U_{*}=\{u_{1},\ldots,u_{q}\}$ uniformly at random from all vertices in $U$ having degree at least $\log\log n$ to $V\setminus U$.
@@ 3.3 Construction of $u$
We construct $u$ in the same manner as in [57, Section 6.3]. Namely, it is constructed implicitly by the following recursion:
@@ 4 Information-theoretic threshold
Part 2 of the theorem — the impossibility result — follows from a percolation style argument. We first show that we may reduce to the case where only the within-community edges are revealed to us. Consider generating a sample from $G\sim\mathsf{SBM}(n,q,d,\lambda)$ as follows. Assign each vertex $v$ an i.i.d. uniform $\sigma^{\star}_{v}\in[q]$ community label as before. For each pair of vertices $u,v$ such that $\sigma^{\star}_{u}=\sigma^{\star}_{v}$ connect them with an edge independently with probability $\frac{qd\lambda}{n}\cdot\frac{1}{1-\frac{d(1-\lambda)}{n}}=(1+o(1))\frac{qd\lambda}{n}$ to obtain a graph $\widetilde{G}$. Finally, let $G=\widetilde{G}\cup\widehat{G}$ where $\widehat{G}$ is the Erdős–Rényi graph $G\left(n,\frac{d(1-\lambda)}{n}\right)$ independent of $\widetilde{G}$. We claim that $\delta$-recovery is possible in $\mathsf{SBM}(n,q,d,\lambda)$ only if recovery is possible using $\widetilde{G}$. Indeed, given a sample of $\widetilde{G}$ we can simply sample $\widehat{G}$ and perform recovery from the block model. Thus, it suffices to show that $\delta$-recovery is impossible from $\widetilde{G}$.